# Thermodynamic uncertainty relations in superconducting junctions

### v1

@@ Abstract
We turn now to the analysis of the SS case, where the presence of MAR processes, which are shown in Fig. 1(e-f), adds additional complexity. Again we refer to Ref. [35, 36, 43] for the current and noise calculation in the framework of the full counting statistics. The critical transmission $\tau_{\rm c}$ for the SS case is shown in Fig. 2(c). It is noticeable that the TUR is broken for generally smaller transmissions in comparison to the NS case [see panel (a)]. For higher temperatures, the critical transmission decreases. In contrast to the NS case, the critical transmission depends drastically on the voltage even for larger temperatures. In particular, it is noticeable that the dependence of the voltage is more pronounced in the vicinity of voltages corresponding to onsets of MAR processes, namely $V=2\Delta/n$, with $n$ being an integer. This suggests that the breaking of TUR is again rooted in the nonlinearity induced by the coexistence of more than one transport process. Also notice that for low temperatures, the TUR violation concentrates on only one point, namely $V=T=0$. This can be explained by considering that in the limit $V,T\to 0$ for an SS junction with high transmission, all transmission probabilities up to infinite order contribute. In the case of perfect transmission, the current does not go to zero for $V,T\to 0$ and converges to a non-zero value [35, 36], rendering the absolute limit of violation, where the TUR-breaking coefficient $\mathcal{F}$ diverges.
@@
The minimal TUR breaking coefficient is displayed in Fig. 2(d), where it is seen that the TUR is severely more broken than in the NS case, reaching absolute values of one order of magnitude higher than in the NS case. This is caused by highly transmissive very high order MAR processes coexisting with low transmissive slighly lower order MAR processes, granting an immensely strong violation of the TUR for a large charge number $n$ in the coefficient $n^{4}p_{n}(1-6p_{n})$. As the voltage range is restricted for $\mathcal{F}_{\rm min}$ in Fig. 2(d) due to computation time, $eV\geq 0.1\Delta$, an even stronger violation is achievable for lower voltages.
@@
Conclusions.– In this work, we have studied the violation of TUR in the coherent charge transport in junctions containing superconductors. We have shown that the presence of multiple tunneling processes tends to lead to violations of TUR and that the degree of violation is larger for higher-order transport processes. We have demonstrated this by providing an analytical formula for the breaking of TUR for small voltages and see that it only depends on the zeroth order transmission probabilities (which are evaluated at equilibrium). In addition, for junctions containing superconductors, the requirement for the energy dependence of the transmission as a source of nonlinearity is not necessary to break TUR. Our work shows that superconducting junctions of simple quantum point contacts are promising candidates for the engineering of quantum thermal machines that can exhibit high efficiencies, being immune against fluctuations in a low dissipation scenario because of their potential to strongly violate TUR.
@@
D.C.O. and W.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation) via SFB 1432 (Project No. 425217212). J.C.C. acknowledges support from the Spanish Ministry of Science and Innovation under Grant No. PID2020-114880GB-I00 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M) and thanks the DFG and SFB 1432 for sponsoring his stay at the University of Konstanz as a Mercator Fellow. R.L acknowledges support by the Spanish State Research Agency (MCIN/AEI/10.13039/501100011033) and FEDER (UE) under Grant No. PID2020-117347GB-I00 and María de Maeztu Project No. CEX2021-001164-M and Grant No. PDR2020/12 sponsored by the Comunitat Autonoma de les Illes Balears through the Direcció General de Política Universitaria i Recerca with funds from the Tourist Stay Tax Law ITS 2017-006. K.W.K. acknowledges the support by the National Research Foundation of Korea (NRF) grant funded by the Korea government(MSIT) (No. 2020R1A5A1016518).

### v2

@@ Abstract
We turn now to the analysis of the SS case, where the presence of MAR processes, which are shown in Fig. 1(e-f), adds additional complexity. Again we refer to Ref. [Cuevas2003, Cuevas2004, Ohnmacht2023] for the current and noise calculation in the framework of the full counting statistics. The critical transmission $\tau_{\rm c}$ for the SS case is shown in Fig. 2(c). It is noticeable that the TUR is broken for generally smaller transmissions in comparison to the NS case [see panel (a)]. For higher temperatures, the critical transmission decreases because the processes overlap more, which benefits the TUR breaking. As compared to the NS case, the critical transmission depends more drastically on the voltage even for higher temperatures. In particular, the dependence of the voltage is more pronounced in the vicinity of voltages corresponding to onsets of MAR processes, namely $V=2\Delta/n$, with $n$ being an integer. Also notice that for low temperatures, the TUR violation concentrates on only one point, namely $V=T=0$. This can be explained by considering that in the limit $V,T\to 0$ for an SS junction with high transmission, all transmission probabilities up to infinite order contribute. In the case of perfect transmission, the current does not go to zero for $V,T\to 0$ and converges to a non-zero value [Cuevas2003, Cuevas2004], rendering the absolute limit of violation.
@@
The minimal TUR breaking coefficient is displayed in Fig. 2(d), where it is seen that TUR is severely more broken than in the NS case, reaching absolute values of one order of magnitude higher than in the NS case. This enhanced breaking must originate from the MARs, which are exclusive to the SS. In particular, this large breaking results from highly probable, high-order MAR processes coexisting with lots of lower probability MARs of lower order, granting an immensely strong violation of the TUR. As the voltage range is restricted for $\mathcal{F}_{\rm min}$ in Fig. 2(d) due to computation time, $V\geq 0.1\Delta$, an even stronger violation is achievable for lower voltages.
@@
Conclusions.– In this work, we have studied the violation of TUR in the coherent charge transport in junctions containing superconductors. We have shown that the presence of multiple tunneling processes help breaking TUR and that the degree of violation is clearly larger in the SS case. We have rationalized this by providing an analytical formula for the TUR-breaking coefficient for low voltages and show that it only depends on the zero-bias tunneling probabilities. We have also shown that for junctions containing superconductors, no energy dependence of the normal transmission coefficient is required to break TUR, which is due to the intrinsic energy dependence of the probability of the different superconducting tunneling processes. This is a key advantage that makes it much more feasible to observe the violation of TUR in superconducting junctions. So, in summary, our work shows that superconducting junctions of simple quantum point contacts are promising candidates for the engineering of quantum thermal machines that are immune against fluctuations in a low dissipation scenario. Our contribution also paves the way for the experimental observation of the TUR breaking in the context of quantum electron transport, something that continues to be very elusive.
@@
D.C.O. and W.B. acknowledge support by the Deutsche Forschungsgemeinschaft (DFG; German Research Foundation) via SFB 1432 (Project No. 425217212). J.C.C. acknowledges support from the Spanish Ministry of Science and Innovation under Grant No. PID2020-114880GB-I00 and through the “María de Maeztu” Programme for Units of Excellence in R&D (CEX2023-001316-M) and thanks the DFG and SFB 1432 for sponsoring his stay at the University of Konstanz as a Mercator Fellow. R.L. acknowledges support by the Spanish State Research Agency (MCIN/AEI/10.13039/501100011033) and FEDER (UE) under Grant No. PID2020-117347GB-I00 and María de Maeztu Project No. CEX2021-001164-M and Grant No. PDR2020/12 sponsored by the Comunitat Autonoma de les Illes Balears through the Direcció General de Política Universitaria i Recerca with funds from the Tourist Stay Tax Law ITS 2017-006. K.W.K. acknowledges the support by the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. 2020R1A5A1016518).